OpenStudy (anonymous):

Solve. x2 – 4x – 6 = 0

OpenStudy (anonymous):

is it x sqr?

OpenStudy (anonymous):

dhashni where r u from?

OpenStudy (anonymous):

me!!!! i am from india

OpenStudy (anonymous):

solve using quadratic formula \[(x = -b \pm \sqrt{b^2 - 4ac})/2a\]

OpenStudy (anonymous):

yes it is, do u want me to draw it for u

OpenStudy (anonymous):

one correction please it is x = ................ /2a

OpenStudy (anonymous):

x = 2+/-sqrt(10)

OpenStudy (anonymous):

x = 5.16, -1.16

OpenStudy (anonymous):

it isnt that simple, i wish it were there are 4 answers given and none of them are what is given here

OpenStudy (nilankshi):

kush write your qus

OpenStudy (anonymous):

so what are the options?

OpenStudy (anonymous):

when you want to show x or any number to the power of some number use = x^2 = \[x^2\]

OpenStudy (anonymous):

x = 2+/-sqrt(10) the only thing is it isnt a negative sqrt it is a positive sqrt

OpenStudy (anonymous):

Possible intermediate steps: x^2-4 x-6 = 0 Add 6 to both sides: x^2-4 x = 6 Add 4 to both sides: x^2-4 x+4 = 10 Factor the left hand side: (x-2)^2 = 10 Take the square root of both sides: abs(x-2) = sqrt(10) Eliminate the absolute value: x-2 = -sqrt(10) or x-2 = sqrt(10) Add 2 to both sides: x = 2-sqrt(10) or x-2 = sqrt(10) Add 2 to both sides: x = 2-sqrt(10) or x = 2+sqrt(10)

OpenStudy (anonymous):

my answers are correct, just not written in the same way. it depends, if you are allowed to use a scientific calculator, or not.

OpenStudy (anonymous):

\[x = 2 -\sqrt{10} \] \[x = \sqrt{10 + 2} \]

OpenStudy (anonymous):

The above are correct. What answers do they give you?

OpenStudy (anonymous):

\[ x \approx -1.1622776601684 \] \[ x \approx 5.1622776601684 \]

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b ^{2}-4ac})/2a\] a=1, b=-4 c=-6 x= \[(-(-4) +\sqrt{(-4)^{2}-4(1)(-6) )}/((2)(1))\] =\[2+\sqrt{5}\] or x= \[(-(-4) -\sqrt{(-4)^{2}-4(1)(-6) )}/((2)(1))\] =\[2-\sqrt{5}\]